2. Valle de Estós

TRACK OF THE RIGAU Recommended to spend half day (morning preferably) and have lunch there

Depart from the Main Square (Plaza Mayor) of Benasque. Cross the Ésera river and continue to your right. The path to be followed is a wide one always on the right of the river. Leave a bridge on your right, cross the Someres ravine a bit later and you will get to the Es Padellases spring. The track goes through Es Sacs, narrows and finally gets to the Ruda Power Station. Turn right and cross a concrete river till you get to a crossroad. If you turn left you will get to the Estós refuges and the Renclusa. If you take it to the right, you will go back to Benàs.